OpenStudy (anonymous):

simplify 2x{x+(x+2)[(x-1)(x+2)-3]+4}-(3x^2+2x-1)

OpenStudy (saifoo.khan):

Somebdy changed her photo!

OpenStudy (anonymous):

One sec editing mistake

OpenStudy (anonymous):

2x{x+(x+2)[(x-1)(x+2)-3]+4}-(3x^2+2x-1) Okay...you have to take this one step at a time. PEMDAS here we go. I am going to look only at: [(x-1)(x+2)-3] for a minute (x^2 + x - 2) - 3 = x^2 + x - 5 now * (x+2) (x+2)(x^2 + x - 5) = x^3 + 3x^2 - 3x - 10 now + x x^3 + 3x^2 - 3x + x - 10 = x^3 + 3x^2 - 2x - 10 now + 4 x^3 + 3x^2 - 2x - 10 + 4 = x^3 + 3x^2 - 2x - 6 Now step back and look at the equation 2x{x+(x+2)[(x-1)(x+2)-3]+4}-(3x^2+2x-1) everything inside the {} is dealt with so multiply by 2x 2x(x^3 + 3x^2 - 2x - 6) = 2x^4 + 6x^3 - 4x^2 - 12x Now let's look subtract! 2x^4 + 6x^3 - 4x^2 - 12x - (3x^2 + 2x - 1) Make sure to watch your signs 2x^4 + 6x^3 - 4x^2 - 12x - 3x^2 - 2x + 1 2x^4 + 6x^3 – 7x^2 – 14x + 1

OpenStudy (anonymous):

There we go

OpenStudy (anonymous):

And yes :D, I changed my pix

OpenStudy (anonymous):

@lizet5, does this make sense??

OpenStudy (anonymous):

fibonacci style @ saifoo

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

yes it does make sense i was just having trouble with the brackets i got most of the answer except for the -14x-1, your answer is the same as my instructor. I got your steps thank you very much

OpenStudy (anonymous):

You are welcome. Sometimes it helps also to write it down with lots of space in between on scratch paper or color code stuff so that it is clear to yourself.

OpenStudy (anonymous):

yeah i have to write on computer paper because the lines confuse me! but i will buy some color pencils that's a great idea. thanks again

OpenStudy (anonymous):

No problem! Good luck